An odd number which when multiplied by 4 and added to the next odd number produces a sum of 177. Find the number.
I got to n+x=44.25. Is this right?

Answer by checkley79(3341)

4X+(X+2)=177
5X+2=177
5X=177-2
5X=175
X=175/5
X=35 ANS. FOR THE FIRS ODD NUMBER.
PROOF:
4*35+(35+2)=177
140+37=177
177=177